OpenStudy (anonymous):

using the appropriate special triangle, 2 sqaureroot 2cosx= 2

OpenStudy (lgbasallote):

and the question is....?

OpenStudy (anonymous):

solve for x ?

OpenStudy (anonymous):

in that case x = 2n(pi) + pi/3

OpenStudy (anonymous):

nono determine if 0 degrees is bigger than 90 degrees or or less

OpenStudy (lgbasallote):

determine if 0 degrees is bigger than 90? of course 90 degrees is bigger than 0! o.O

OpenStudy (anonymous):

nono like yeah I guess solve, and find out how many degrees it is

OpenStudy (anonymous):

... make up your mind plz

OpenStudy (anonymous):

:D !!

OpenStudy (anonymous):

I have to solve the question ... using special triangles

OpenStudy (lgbasallote):

....so it is solve for x....?

OpenStudy (anonymous):

try this way: 2\[\sqrt{2cosx}\]=2....divide both sides by 2 u'll get \[\sqrt{2cosx}\]=1.....now square both sides u'll get 2cosx=1..............divide both sides by 2 u'll get cosx=1/2......can u figure out the rest from here?